You are considering buying a call option for Bernie Bros, a company that is building a series of assisted living homes for aging hippies. Bernie Bros current stock price is $42.20, and the call option you are looking at sells for $5.85 with a $40.00 strike price and six months to expiration.

a. What is the intrinsic value of this option today?

b. What is the premium of this option today?

c. Draw a payoff graph for this option with the stock price at expiration on the x-axis. Plot the payoffs for both the buyer and the seller of this call option. d. If the perceived future volatility of the shares of Bernie Bros was to decrease, what would happen to the price of this option?

e. If you decided you wanted to buy a one-year option instead of the six months, would it cost you more or less? Why?
Answer #1

.a .Intrinsic value is the amount of gain that can be made by immediately exercising the option.

Strike price of call option =$40.00

The option gives the buyer the right to buy at $40.00

The current market price =$42.20

The buyer, if he is allowed to exercise the option by buying a share at $40,he can sell it in the market for $42.20and make a gain of (42.20-40.00)=$2.20

Hence, intrinsic value of the option =$2.20

Intrinsic value of a call option =Current Market Price- Strike Price

.b. Option premium is the price paid to acquire the OPTION.

Option premium=$5.85

.c. Payoff Diagram.

Strike Price =$40.00

Assume Price at expiration =S

If S < or=$40.00,

Payoff for Long call (Buy call)=$0

Payoff for Putcall (SELL call)=$0

If S>$40.00,

Payoff for Long call (Buy call)=(S-40.00)

Payoff for Put call (SELL call)=(40.00-S)

S

A=S-40 if S>40; otherwise=0

B=40-S,if S>40:otherwise =0

Price at Expiration

Payoff for CALL Buyer(Long)

Payoff for CALL Seller(Short)

$30

$0

$0

$31

$0

$0

$32

$0

$0

$33

$0

$0

$34

$0

$0

$35

$0

$0

$36

$0

$0

$37

$0

$0

$38

$0

$0

$39

$0

$0

$40

$0

$0

$41

$1

($1)

$42

$2

($2)

$43

$3

($3)

$44

$4

($4)

$45

$5

($5)

$46

$6

($6)

$47

$7

($7)

$48

$8

($8)

$49

$9

($9)

$50

$10

($10)

.d. If volatility is decreased, Option premium will also decrease

.e. One year option will cost more.

Option Premium=Intrinsic Value+ Time Value

If it is increased to one year from six months, Tine Value will increase